uppose that the duration of a particular type of criminal trial is known to have a mean of 21 days and a standard deviation of 7 days. We randomly sample 9 trials. X ~ n Correct: Your answer is correct. 189 Correct: Your answer is correct. , 21 Correct: Your answer is correct. Part (c) Find the probability that the total length of the 9 trials is at least 214 days. (Round your answer to four decimal places.) Part (d) Seventy percent of the total of 9 of these types of trials will last at least how long? (Round your answer to two decimal places.) daysX ~ n Correct: Your answer is correct. 189 Correct: Your answer is correct. , 21 Correct: Your answer is correct. Part (c) Find the probability that the total length of the 9 trials is at least 214 days. (Round your answer to four decimal places.) Part (d) Seventy percent of the total of 9 of these types of trials will last at least how long? (Round your answer to two decimal places.) days
Answer #1

c) probability that the total length of the 9 trials is at least 214 days =P(X>214)=P(Z>(214-189)/21)=P(Z>1.19)=0.8830

(try 0.1169 if z score is required till 3 decimal places)

d)

for 30th percentile critical value of z= -0.52
therefore corresponding value=mean+z*std deviation= 178.08

(try 178.00 if this comes wrong for f z score is required till 3 decimal places)